# Odd coloring graphs with linear neighborhood complexity

### v1

@@ 5 Odd coloring and bipartite circle graphs
Finally, we use one set of $\ell$ palettes of $9$ colors to color $B_{i}$ for even $i$, and a second set of $\ell$ palettes of $9$ colors to color $B_{i}$ for odd $i$. By Claim 3, this results in a coloring $\phi$ with $2\cdot 9\cdot\ell$ colors that is odd for $\{C^{*}(T,f)\colon f\in T\}$. Using further, that by Lemma 5.1 in none of the palettes the color $9$ is used as the odd color in a neighborhood, we can merge all these, yielding a total of $12\cdot 8\cdot\ell+$ 1 colors. ∎
We now use Theorem 5.2 to prove Corollary 1.4, which is restated below.
See 1.4
@@
De Fraysseix [16] proved that every connected bipartite circle graph is a fundamental graph of a planar graph $G$ and a spanning tree $T$. So $G$ excludes $K_{5}$ as a minor, and the graphs excluding $K_{4}$ as a minor are exactly the graphs of treewidth at most $2$. Hence, by [17] they have star arboricity at most $3$. Thus, we can use Theorem 5.2 to color $E(G)\setminus E(T)$ with $16\cdot 3+1=49$ colors so that for each non-bridge edge $f\in E(T)$, there exists a color $i\in\{1,2,\ldots,49\}$ which occurs an odd number of times on the fundamental cut of $f$.
Now, by planar duality $M(G^{*})\cong M(G)^{*}$, see e.g., [43, Chapter 6.3, Corollary 1]. Thus, by Lemma 2.2 the fundamental graph of $G$ and $T$ is isomorphic to the fundamental of graph $G^{*}$ with spanning forest $T^{*}=E(G)\setminus E(T)$. Applying Theorem 5.2 to this graph, we can use a second set of size $49$ to color $E(G^{*})\setminus E(T^{*})=E(T)$ such that for each non-bridge edge $f\in E(T^{*})=E(G)\setminus E(T)$, there exists a color $i\in\{50,51,\ldots,98\}$ which occurs an odd number of times on the fundamental cut of $f$ in $G^{*}$ with respect to $T^{*}$. ∎
@@
Does every bipartite circle graph have odd chromatic number at most $4$?
We think that Theorem 5.2 is interesting in its own right and in fact we are not aware of any family of graphs that forces an unbounded number of colors when coloring non-tree edges as in the theorem. In other words, we wonder if the following holds.
Problem 5.4.
Does there exist a constant $k$ so that if $G$ is a graph and $T$ is a spanning forest of $G$, then there exists a coloring $\phi:E(G)\setminus E(T)\rightarrow\{1,\ldots,k\}$ so that for each non-bridge edge $f\in E(T)$, there exists a color $i\in\{1,\ldots,k\}$ which occurs an odd number of times on the fundamental cut $C^{*}(T,f)$?

### v2

@@ 5 Odd coloring fundamental graphs of graphs excluding a minor
Finally, we use one set of $\ell$ palettes of $9$ colors to color $B_{i}$ for even $i$, and a second set of $\ell$ palettes of $9$ colors to color $B_{i}$ for odd $i$. By Claim˜3, this results in a coloring $\phi$ with $2\cdot 9\cdot\ell$ colors that is odd for $\{C^{*}(T,f)\colon f\in T\}$. Using further, that by Lemma˜5.1 in none of the palettes the color $9$ is used as the odd color in a neighborhood, we can merge all these, yielding a total of $2\cdot 8\cdot\ell+1$ colors. ∎
We now use Theorem˜1.5 to prove Corollary˜1.6, which is restated below.
See 1.6
@@
De Fraysseix [27] proved that every connected bipartite circle graph is a fundamental graph of a planar graph $G$ and a spanning tree $T$. So $G$ excludes $K_{5}$ as a minor, and the graphs excluding $K_{4}$ as a minor are exactly the graphs of treewidth at most $2$. Hence, by [21] they have star-arboricity at most $3$. Thus, we can use Theorem˜1.5 to color $E(G)\setminus E(T)$ with $16\cdot 3+1=49$ colors so that for each non-bridge edge $f\in E(T)$, there exists a color $i\in\{1,2,\ldots,49\}$ which occurs an odd number of times on the fundamental cut of $f$.
Now, if $G^{*}$ denotes the planar dual graph of $G$, the we have $M(G^{*})\cong M(G)^{*}$, see e.g., [59, Chapter 6.3, Corollary 1]. Thus, by Lemma˜2.2 the fundamental graph of $G$ and $T$ is isomorphic to the fundamental of graph $G^{*}$ with spanning forest $T^{*}=E(G)\setminus E(T)$. Applying Theorem˜1.5 to this graph, we can use a second set of size $49$ to color $E(G^{*})\setminus E(T^{*})=E(T)$ such that for each non-bridge edge $f\in E(T^{*})=E(G)\setminus E(T)$, there exists a color $i\in\{50,51,\ldots,98\}$ which occurs an odd number of times on the fundamental cut of $f$ in $G^{*}$ with respect to $T^{*}$. ∎
See 1.7
Proof.
Let $G_{n}=(V,E)$ be the circle graph given by vertex set $V=\{(a,b)\mid 1\leqslant a<b\leqslant n\}$ and edge set $E=\{\{(a,b),(c,d)\}\mid a<c<b<d\text{ cyclically}\}$, see Figure˜1 for an example.
Figure 1: A representation of $G_{6}$ as circle graph.
Suppose now that $f:V\to[k]$ is an (improper) odd $k$-coloring and denote by $o_{f}:V\to 2^{[k]}\setminus\{\emptyset\}$ the function that to every vertex of $G_{n}$ assigns the (non-empty) set of colors that appear an odd number of times in the neighborhood of the vertex. If $n\geqslant 2^{k}$ then there are vertices $(a,b),(a,c)$ with $o_{f}(a,b)=o_{f}(a,c)$. But now, observe that $N(b,c)=N(a,b)\Delta N(a,c)$ and hence $o_{f}(b,c)=o_{f}(a,b)\Delta o_{f}(a,c)=\emptyset$, which is a contradiction. ∎
@@
Problem 5.2.
Does every bipartite circle graph have odd chromatic number at most $4$?
We think that Theorem˜1.5 is interesting in its own right and in fact we are not aware of any family of graphs that forces an unbounded number of colors when coloring non-tree edges as in the theorem. In other words, we wonder if the following holds.
@@
Does there exist a constant $k$ so that if $T$ is a spanning forest of a graph $G$, then there exists a coloring $\phi:E(G)\setminus E(T)\rightarrow\{1,\ldots,k\}$ so that for each non-bridge $f\in E(T)$, there exists a color $i\in\{1,\ldots,k\}$ which occurs an odd number of times on the fundamental cut of $T$ with respect to $f$?